OpenStudy (anonymous):

Verify that sin(x^(1/3)) is an odd function and use that fact to show that...

OpenStudy (anonymous):

Verify that \[f(x)=\sin \sqrt[3]{x}\] is an odd function and use that fact to show that \[0\le \int\limits_{-2}^{3}\sin \sqrt[3]{x}dx \le1\] I know how to verify of the function is even or odd, but I'm having trouble showing why the inequality is true.

OpenStudy (anonymous):

Have you ever tried the mean value theorem??

OpenStudy (anonymous):

Hmm. I hadn't thought of that. I can try that and see how it goes.

OpenStudy (anonymous):

Actually, I don't know if I can do that. I need to find the antiderivative of the function in order to use the Mean Value Theorem and I'm not even sure how to find the antiderivative of this function.

OpenStudy (anonymous):

as far as I know you don't have to find the antiderivative in the mean value theorem! It says that